# BED-LLM: Intelligent Information Gathering with LLMs and Bayesian Experimental Design

### v1

@@ C.2 Generating candidate hypotheses
Figure 7: Prompt for generating hypotheses (and evaluating their probability) for the data–estimation method.
@@
Figure 8: An example of the sample distribution generated using the prompt in Figure˜7, conditioned on the two question/answer pairs at the top of this figure. At this stage of the game, we independently sample 200 hypotheses from the LLM and record their frequencies. Note that the distribution exhibits strong mode collapse, with most of the mass highly concentrated on just a few answers, which negatively impacts the performance of the data–estimation method. This summary is for diagnostic purposes: Algorithm˜1 operates on the probabilities of individual samples and never instantiates such an aggregated summary.
Appendix D Experimental details
D.1 Generation of the Animals Dataset
@@
To create the Animals dataset, we prompted OpenAI o3 (OpenAI,, 2025, o3-2025-04-16) to generate a list of animals, using the prompt in Figure˜9. The resulting list is shown in Figure˜10. Alternative names (after $|$ character) were manually added.
@@
You are a zoologist. Please generate a list of 100 living animal species with very high taxonomic diversity, including diversity in phyla, classes, orders, and families. Present each animal on a different line.
Figure 9: Prompt for Animals dataset generation.
@@
Figure 10: Animals dataset (generated using OpenAI o3, with manual curation)
D.2 LLM models details
The open-weights LLM models used in Section˜6 were obtained from HuggingFace (HF) (Wolf et al,, 2020) and their repositories are: mistralai/Mistral-Large-Instruct-2411, Qwen/Qwen2.5-72B-Instruct, meta-llama/Llama-3.3-70B-Instruct. Since all models are instruction-finetuned, we used the latest chat templates provided as part of the HF implementation. All of the experiments using mistralai/Mistral-Large-Instruct-2411 were conducted by Oxford researchers on Oxford compute infrastructure.
The OpenAI API models used were: gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06.
Appendix E Additional Personalization Experiments
E.1 EIG estimates of questions chosen by different questioners
Figure 11 demonstrates the progression of EIG estimates throughout the turns for BED-LLM and Naive QA. We see that at turn $t=1$, the questions chosen by both methods have similar EIG estimates, which explains the similar performance of BED-LLM and Naive QA at turn $t=1$.
Figure 11: The average information gains for each strategy at each turn for the Qwen2.5-72B vs Qwen2.5-72B example. We see that the initial questions have similar information gains, which explains why the performance of the Naive QA method outperforms EIG at turn 1. Afterwards, the EIG information gains are much larger while the Baseline information gains decrease sharply.

### v2

@@ D.2 Generating candidate hypotheses
Figure˜5 shows an example of the distribution of samples obtained following two rounds in the 20–questions game. Note that the samples are highly concentrated on just a handful of answers. This lack of diversity shows that the model’s belief distribution is far more concentrated relative to the variability over valid hypotheses in the ground–truth task distribution, which negatively impacts the performance of the data–estimation method.
@@
Figure 5: An example of the sample distribution generated using the prompt in Section˜D.2, conditioned on the two question/answer pairs at the top of this figure. At this stage of the game, we independently sample 200 hypotheses from the LLM and record their frequencies. Note that the distribution exhibits strong mode collapse, with most of the mass highly concentrated on just a few answers, which negatively impacts the performance of the data–estimation method. This summary is for diagnostic purposes: Algorithm˜1 operates on the probabilities of individual samples and never instantiates such an aggregated summary.
Appendix E Generating candidate hypotheses for BED-LLM
Figure 6: Prompt for generating candidate hypotheses for the “Things” dataset. Similar prompts were used for “Celebrities” and “Animals”. You are playing a game of 20 Questions. Using all of the questions and answers so far:
Generate up to {num_samples} candidate entities that satisfy every clue.
Each candidate must be a single, self-contained entity (e.g., "Europa", "Bagpipe", "Diadem").
List each entity on its own line - no numbering, punctuation, or extra text.
Produce a varied set by identifying features not implied by the clues and diversifying along them.
Do not repeat any entity.
Return only the list of entities.
A fundamental challenge for BED-LLM and its ablations is generating a sufficiently diverse set of candidate hypotheses from the LLM’s belief distribution, that are consistent with the previously-answered questions. Below, we detail the steps we take to construct our distribution over hypotheses.
Candidate hypotheses are generated jointly, rather than independently.
[…]
$p_{\mathrm{LLM}}(\theta^{(1)}_{t},\ldots,\theta^{(N)}_{t};h_{t})\;=\;\prod_{n=1}^{N}p_{\mathrm{LLM}}(\theta_{t}^{(n)};[\theta_{t}^{(1:n-1)},h_{t}])\,.$
We use a diversity-encouraging prompt
We use a prompt designed to elicit stratified hypotheses by encouraging the LLM to consider different semantic features (e.g. age groups, genres, or categories) and implicitly diversify across them. An example prompt is shown in Appendix˜E. In our generation prompt, we reverse the order of the question–answer pairs in $h_{t}$ to place the most recent question at the top of the context window (while retaining earlier exchanges), ensuring that specific constraints are prioritized and mitigating context drift. For the 20 Questions experiments, we used a higher-than-normal temperature ($T=1.3$) to increase diversity of responses. For the preference elicitation experiments, we used $T=1$ to obtain more coherent responses.
Candidates are filtered based on the history
For each candidate, we use $p_{\mathrm{LLM}}(\theta;h_{t-1})$ to assess whether it is compatible with the previous question/answer pairs. We filter responses where the probability of the given answer falls below a certain threshold; in our experiments we set this threshold to 0.2.
Valid candidates from previous generations are included
We also filter the candidate hypotheses from the previous generation, based on the most recent question/answer pair, and include these in our candidate set. We repeat the generation process, keeping the previously generated and filtered samples in context to elicit new generations, either twice or three times if sufficient hypotheses have not been generated (noting the number of possible valid samples can be less than the number requested).
We assume a uniform distribution over hypotheses
While one could in principle reweight candidates using importance sampling, in practice we choose to not rely on the model’s internal probabilities. Instead we approximate the prior as a uniform distribution over this union
$p_{f}(\theta;h_{t})\approx\frac{1}{|\Theta|}\sum_{\theta\in\Theta}\delta_{\theta}.$
Finally, we note that different LLMs respond differently to strategies aiming to increase diversity: some benefit more from a higher temperature while others benefit from more repetitions of the sampling–filtering cycle. For fairness, in our experiments we have kept these parameters constant across models.
Appendix F Experiment details for 20 Questions
F.1 Problem sets
We evaluate across three distinct problem sets—Animals, Celebrities, or Things—with each containing a mix of 100 obscure and common targets. Here, the problem set is just a list of different $\theta^{*}$ that will be individually provided to the answerer to instantiate different problems (e.g. we conduct a trial where $\theta^{*}=$“dog”, then one where $\theta^{*}=$“cat”, etc). The list of targets is never provided to the questioner model to restrict the set of possible hypotheses: the questioner is only prompted that is trying to identify an “animal”, “celebrity”, or “thing”. The problem sets are
  • Animals: a set of animal species generated with OpenAI’s o3 model to ensure a diverse mix and balanced taxonomy.
  • Celebrities: a diverse set of public figures, as used by Zhang et al. (2024).
  • Things: a collection of everyday and exotic entities drawn from the web corpus, as used by Zhang et al. (2024); it covers a wide range of categories, from plants and clothing to professions, events, and mythical creatures.
@@ F.1 Problem sets
To create the Animals problem set, we prompted OpenAI o3 (OpenAI, 2025, o3-2025-04-16) to generate a list of animals, using the prompt in Section˜F.1. The resulting list is shown in Figure˜8. Alternative names (after $|$ character) were manually added.
Figure 7: Prompt for Animals problem set generation. You are a zoologist.
@@
Please generate a list of 100 living animal species with very high taxonomic diversity, including diversity in phyla, classes, orders, and families. Present each animal on a different line.
@@
Figure 8: Animals problem set (generated using OpenAI o3, with manual curation)
F.2 Evaluation
[…]
Figure 9: Full plots for 20 Questions Experiments.
F.3 Algorithmic details
Using our sample–then–filter process (see Section˜3.3), we aim to sample at least $N=15$ hypotheses, repeating the cycle up to three times if needed (the exact number of hypotheses can be less than this as it may not be possible to generate sufficient valid hypotheses, especially in later experiment turns). The questioner generates $M=15$ candidate questions to test, $\mathcal{X}^{\text{cand}}$, using the “conditional generation” approach of Section˜3.1 when possible, but falling back on “unconditional generation” if insufficient candidate hypotheses have been generated.
Appendix G Experiment details for preference elicitation
G.1 Problems
To generate a set of ground-truth user profiles, we take a set of $200$ real user ratings from the MovieLens-100K dataset (Harper & Konstan, 2015), then use an “oracle” LLM (namely, OpenAI’s o3 model) to produce a paragraph of text that is consistent with each distinct set of ratings, using the prompt in Section˜G.3. As was the case for the 20 Questions problems, this problem set is never provided to the questioner and the set of allowed $\theta$ is not constrained.
Because we need the LLM to be able to meaningfully capture uncertainty in the space of responses, we restrict questions to be multiple-choice. Specifically, the questioner is tasked with producing a question along with five possible responses A/B/C/D/E. We then define each $x_{t}$ to be the question coupled with the possible responses, and each $y_{t}$ to be one of the letters A/B/C/D/E to provide a restricted set of tokens over which we can measure entropy. Option E is further constrained to always be “none of the above” so that the answerer is not committed to choosing one of the directly generated choices if none are suitable.
G.2 Evaluation
To again allow tracking of progress through the experiment, after each turn of the interaction $t$, the questioner generates ten film recommendations, conditioned on $h_{t-1}$. These recommendations are checked for consistency with prior questions and answers; if any are judged inconsistent then they are removed and additional recommendations are generated. The quality of the film recommendations is then assessed using an “LLM-as-judge” protocol (Zhu et al., 2025; Trivedi et al., 2024). Namely, the answerer evaluates each of the 10 films recommended by the questioner, conditioned on the hidden target user profile $\theta^{*}$. It scores each film on a scale of 1 to 5 (in 0.5 increments), based on how well the recommendation aligns with $\theta^{*}$ — this score is output together with a brief justification to increase reliability. We report the mean rating and standard error across 200 users, over 5 question–answer turns.
G.3 Algorithmic details
For BED-LLM and Entropy, we compare $M=8$ candidate questions at each turn and we aim to generate at least $N=5$ candidate hypotheses. We use the “unconstrained generation” approach of candidate question generation (see Section˜3.1) as the user profiles can be quite diffuse and we are only generating a small number of possible hypotheses that can be quite easy to split.
We note that data–estimation setup is not at all viable for this problem because the large number of tokens and varying dimensionality of each $\theta$ sample mean that $\mathrm{H}[p_{\mathrm{LLM}}(\theta;[h_{t-1},x_{t+1},y_{t+1}])]$ is not only infeasible to estimate, but also is not a meaningful measure of uncertainty.
Figure 10: Prompt used to generate ground-truth user profiles for preference elicitation task. You will be given a user’s complete film rating history from the MovieLens dataset, provided as a dictionary structured by rating levels.
Your task is to thoroughly analyze the user’s preferences across the entire range of their film ratings (from highest to lowest). Then, write a cohesive, descriptive paragraph (approximately 5–7 sentences) summarizing the user’s overall film taste profile.
In your response, explicitly address:
Favored Elements (inferred primarily from 4.5–5.0 ratings): Highlight the genres, narrative styles, themes, tones, historical eras, and emotional experiences that consistently resonate positively with this user. Avoid mentioning any specific film titles, characters, or explicit plot points. Neutral or Mixed Preferences (inferred from ratings around 2.5–4.0): Note if there are indications of genre overlap or conditional enjoyment, such as certain genres or styles they occasionally appreciate under specific circumstances. Disliked Elements (inferred primarily from 0.5–1.5 ratings): Clearly outline the genres, narrative characteristics, tones, or emotional impacts that the user consistently finds unappealing or poorly executed. Your paragraph must be precise, informative, nuanced, and balanced, effectively capturing the complexity and specificity of the user’s movie preferences. The resulting profile should be clear and detailed enough for a recommendation system to accurately predict the user’s likely enjoyment or dislike of other films based on their established patterns of taste.
Proceed carefully, reasoning explicitly about the user’s overall rating patterns rather than relying exclusively on extreme ratings, to form a comprehensive, stable, and representative film preference profile.